Solve the equation by factoring and show the complete step by step solution.
x^2-2x+2=0

Answer by Theo(13342)

this doesn't factor easily, so you need to use the quadratic formula.

that formula is:



               -b plus or minus square root of (b^2 - 4ac)
     x =     -----------------------------------------------
                               2a


the equation needs to be put in standard form of ax^2 + bx + c = 0

it's already in standard form.

this means that a = 1, b = -2, c = 2

plug those values into the quadratic formula and you get:



               2 plus or minus square root of (4 - 4*1*2))
     x =     -----------------------------------------------
                               2*1


you get:



               2 plus or minus square root of (-4)
     x =     -----------------------------------------------
                               2


square root of (-4) is equal to square root of (4 * (-1)) which is equal to 2 * square root of (-1).

square root of (-1) is equal to the letter i.

therefore square root of (-4) is equal to 2*i.

the value of the equation is equal to 0 when:

x = (2 + 2*i)/2 or x = (2 - 2*i)/2

simplify this to get:

x = 1 + i or x = 1 - i

use of a quadratic equation solver online confirms the solution is correct, as shown below:

$$$